OpenStudy (anonymous):

find linear equation whosegraph is the straight line with the given property trough (1, -7/8) with slope (1/8).. please help

OpenStudy (anonymous):

\[y-y_1=m(x-x_1)\] \[y-(-\frac{7}{8})=\frac{1}{8}(x-1)\] \[y+\frac{7}{8}=\frac{1}{8}x-\frac{1}{8}\] \[y=\frac{1}{8}x-1\]

OpenStudy (anonymous):

thank u....